# SecBench: A Comprehensive Multi-Dimensional Benchmarking Dataset for LLMs in Cybersecurity

### v2

@@ Abstract.
Evaluating Large Language Models (LLMs) is crucial for understanding their capabilities and limitations across various applications, including natural language processing and code generation. Existing benchmarks like MMLU, C-Eval, and HumanEval assess general LLM performance but lack focus on specific expert domains such as cybersecurity. Previous attempts to create cybersecurity datasets have faced limitations, including insufficient data volume and a reliance on multiple-choice questions (MCQs). To address these gaps, we propose SecBench, a multi-dimensional benchmarking dataset designed to evaluate LLMs in the cybersecurity domain. SecBench includes questions in various formats (MCQs and short-answer questions (SAQs)), at different capability levels (Knowledge Retention and Logical Reasoning), in multiple languages (Chinese and English), and across various sub-domains. The dataset was constructed by collecting high-quality data from open sources and organizing a Cybersecurity Question Design Contest, resulting in 44,823 MCQs and 3,087 SAQs. Particularly, we used the powerful while cost-effective LLMs to (1). label the data and (2). constructing a grading agent for automatic evaluation of SAQs. Benchmarking results on 13 SOTA LLMs demonstrate the usability of SecBench, which is arguably the largest and most comprehensive benchmark dataset for LLMs in cybersecurity. More information about SecBench can be found at our website (sec, 2024), and the dataset can be accessed via the artifact link (Sec, 2024).
@@ 1. Introduction
To achieve automatic evaluation of SAQs, we employed another powerful yet cost-efficient LLM, GPT-4o-mini, as a grading agent to automatically grade the tested LLMs’ answers based on the question stem and ground truth (correct answer) provided by SecBench. For evaluation, we conducted benchmarking of 13 State-of-the-Art LLMs on SecBench, demonstrating the usability of SecBench both qualitatively and quantitatively.
@@ 5. Evaluation
Based on SecBench, we conducted extensive benchmarking on 13 SOTA LLMs, including the GPT series and competitive open-source ones.
@@ 5.2. SAQ Benchmarking
Table 2 presents the benchmarking results for the 3,087 short-answer questions (SAQs). The values in each cell represent the average score, graded by the grading agent, on a percentage scale for the corresponding category. Compared to MCQs, a larger gap is observed between different LLMs, indicating that solving SAQs is more challenging than MCQs. This is because the tested LLMs are required to generate their own answers rather than simply choosing from given options. Notably, for SAQs, the Tencent Hunyuan-Turbo model (hun, 2024) also outperforms most existing models, achieving a score of 82.13. It ranks second only to the state-of-the-art GPT-4o (85.17) and is competitive with the GPT-4o-mini (82.49).

### v3

@@ Abstract.
Evaluating Large Language Models (LLMs) is crucial for understanding their capabilities and limitations across various applications, including natural language processing and code generation. Existing benchmarks like MMLU, C-Eval, and HumanEval assess general LLM performance but lack focus on specific expert domains such as cybersecurity. Previous attempts to create cybersecurity datasets have faced limitations, including insufficient data volume and a reliance on multiple-choice questions (MCQs). To address these gaps, we propose SecBench, a multi-dimensional benchmarking dataset designed to evaluate LLMs in the cybersecurity domain. SecBench includes questions in various formats (MCQs and short-answer questions (SAQs)), at different capability levels (Knowledge Retention and Logical Reasoning), in multiple languages (Chinese and English), and across various sub-domains. The dataset was constructed by collecting high-quality data from open sources and organizing a Cybersecurity Question Design Contest, resulting in 44,823 MCQs and 3,087 SAQs. Particularly, we used the powerful while cost-effective LLMs to (1). label the data and (2). constructing a grading agent for automatic evaluation of SAQs. Benchmarking results on 16 SOTA LLMs demonstrate the usability of SecBench, which is arguably the largest and most comprehensive benchmark dataset for LLMs in cybersecurity. More information about SecBench can be found at our website (sec, 2024), and the dataset can be accessed via the artifact link (Sec, 2024).
@@ 1. Introduction
To achieve automatic evaluation of SAQs, we employed another powerful yet cost-efficient LLM, GPT-4o-mini, as a grading agent to automatically grade the tested LLMs’ answers based on the question stem and ground truth (correct answer) provided by SecBench. For evaluation, we conducted benchmarking of 16 State-of-the-Art LLMs on SecBench, demonstrating the usability of SecBench both qualitatively and quantitatively.
@@ 5. Evaluation
Based on SecBench, we conducted extensive benchmarking on 16 SOTA LLMs, including the GPT series and competitive open-source ones.
@@ 5.2. SAQ Benchmarking
Table 2 presents the benchmarking results for the 3,087 short-answer questions (SAQs). The values in each cell represent the average score, graded by the grading agent, on a percentage scale for the corresponding category. Compared to MCQs, a larger gap is observed between different LLMs, indicating that solving SAQs is more challenging than MCQs. This is because the tested LLMs are required to generate their own answers rather than simply choosing from given options. Specifically, for SAQs, the o1-preview (gpt, 2024f) and o1-mini (gpt, 2024e) achieved the highest scores (89.24 and 87.50, respectively), due to their advanced capability on reasoning. Note that Tencent Hunyuan-Turbo model (hun, 2024) also outperforms most existing models, achieving a high score of 82.13, which is competitive with the GPT-4o-mini (gpt, 2024c) (82.49) and DeepSeek-V3 (dee, 2024b) (83.71).